Assume you would need 5 RNs 24/7 (all 3 shifts) for your unit/department and the rotation would be every other weekend. Predict how many worked FTEs would be required to adequately staff the unit. Point out how you arrived at your answer.
Answer #1

For adequately staff the unit FTE required are

Hours in a year = 365*24= 8760

So this much hour is required time.

An full time employee can work for 40 hours .

Here no. Of RNs is 5

So work hours will be 40*5*52= 10400

So required FTE is = 10400/8760= 1.18 is required no FTE daily.